OpenStudy (anonymous):

Will someone please help me use L'hopital to evaluate the limit as x-->0+ x^x-1/lnx+x-1

OpenStudy (anonymous):

\[\lim_{x \rightarrow o+} x^x-1/lnx+x-1\]

OpenStudy (anonymous):

I can't see how this would be or could be re-written in the indeterminate form...

OpenStudy (dumbcow):

hmm i don't see it either ... i know the limit is 0 though

OpenStudy (anonymous):

how do you know it is zero?

OpenStudy (dumbcow):

plug in consecutive small values for x ... it seems to go toward 0 also http://www.wolframalpha.com/input/?i=lim+%28x^x-1%29%2F%28log%28x%29%2Bx-1%29+as+x-%3E0%2B :)

OpenStudy (anonymous):

i am not sure what the expression is is it \[\frac{x^x-1}{\ln(x)+x-1}\]?

OpenStudy (anonymous):

if so you cannot use l'hopital because it is not in the form \(\frac{0}{0}\) the form would be \(\frac{-1}{-\infty}\)

OpenStudy (anonymous):

do I rewrite it then? or solve another way?

OpenStudy (anonymous):

it is zero

OpenStudy (anonymous):

no work, just your eyeballs numerator approaches -1, denominator goes to \(-\infty\)

OpenStudy (anonymous):

so the limit as x approaches 0+ of x^x-1 is -1 because as x^x is basically zero. And therefore x^x -1 is -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ooooh no no no sorry that is wrong!!

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

thank you. Our class moves so fast I don't always absorb stuff well...that makes sense....and -1 over infinity is essentially zero

OpenStudy (anonymous):

\[\lim_{x\to 0}x^x=1\]

OpenStudy (anonymous):

i made a mistake sorry

OpenStudy (anonymous):

the limit in the numerator is \(0\) not \(-1\)

OpenStudy (anonymous):

so I have to re write the equation, so I can use L'hopitals rule?

OpenStudy (anonymous):

but that doesn't make any difference to the solution, because as \(x\to 0\) we have \[\ln(x)\to -\infty\]

OpenStudy (anonymous):

no you cannot use l'hopital, because the denominator does not go to zero

OpenStudy (anonymous):

if it was \(\frac{0}{0}\) you could use l'hopital, but it is not

OpenStudy (anonymous):

the numerator goes to zero the denominator goes to \(-\infty\)

OpenStudy (anonymous):

so how do I evaluate this?

OpenStudy (anonymous):

\(\frac{0}{\infty}=0\) it is not indeterminate

OpenStudy (anonymous):

you have a fraction the numerator is getting closer and closer to zero, but the denominator is going to \(-\infty\)

OpenStudy (anonymous):

ok...again..you make sense...so it is just zero then for the answer?

OpenStudy (anonymous):

the ratio is getting smaller very fast

OpenStudy (anonymous):

yes, zero is the answer if the denominator was also going to zero, you would have more work to do, but since the denominator is going to minus infinity, you have nothing to do

OpenStudy (anonymous):

oh ok...I think I almost get it...the zero isn't really a zero...it is just essentially a zero because it is a number getting more and more tiny...so a number over infinity is essentially zero, which is the answer????

OpenStudy (anonymous):

i think i see the confusion \[\frac{0}{\infty}=0\] but also \(\frac{a}{\infty}=0\)

OpenStudy (anonymous):

they are both zero in the limit if the numerator goes to a finite number, and the denominator goes to infinity, then the limit is zero

OpenStudy (anonymous):

for example \[\lim_{x\to 0^+}\frac{100}{\ln(x)}=0\]

OpenStudy (anonymous):

got it...thank you :)